A college student pays $280 for a three hour course. How much would he have to pay for a class load of 16 hours? Round your answer to the nearest whole dollar. Be sure to put a dollar sign ($) on your answer.__________
: A college student pays $280 for a three hour course. How much would he have to pay for a class load of 16 hours? Round your answer to the nearest whole dollar. Be sure to put a dollar sign ($) on your answer.__________

Answer by edjones(2401)

x/16=280/3
16x/16=16*280/3
x=$1493.
.
Ed